OpenStudy (anonymous):

How to write this into equation with a standard form with integer coefficient. y= 3/5x+3

OpenStudy (amistre64):

well, Ax + By = C is standard form, so lets math over the x parts and multiply as needed

OpenStudy (anonymous):

We'll I just got to school and gotta catch up to this, I'm thinking add 3/5x to each side and combine like terms?

OpenStudy (amistre64):

subtract, but yeah

OpenStudy (amistre64):

y= 3/5x+3 -3/5x + y = 3 now we can multiply thru by -5 to shape this up

OpenStudy (anonymous):

How do you multiply thru???

OpenStudy (amistre64):

distribute ... multiply everything in there by -5

OpenStudy (amistre64):

-5 [-3/5x + y = 3] -3(-5)/5 x + (-5)y = 3(-5)

OpenStudy (amistre64):

In standard form the A part of Ax+By =C is not negative,which is why multiplying thru by -5 is required

OpenStudy (anonymous):

Ahh I'm sorry for late response, my connection is low since I'm on iPad anyways is the answer 3x+y=3

OpenStudy (amistre64):

i get 3x -5y = -15

OpenStudy (amistre64):

the site is not working right at the moment so notifications arent geting sent and you have to refresh the page